OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18

OpenStudy (anonymous):

Answers: 12x + 4y = -12 15x = -30 8x + 4y = -4 14x = -10 6x - 8y = -36 -6y = -42 6x - y = -15 3y = 9

OpenStudy (cwrw238):

solve each of the 4 options for x and y and see if they fit the equations in the question for example for first option x = -30/15 = -2 now find y by plugging x=-2 in first equation and see if the results fit

OpenStudy (anonymous):

OKay well im still a little confused.. IM not to sure how to solve these in the first place.. could you walk me through one?

OpenStudy (cwrw238):

I'll find y for you in the first option: 12x + 4y = -12 now we found x = -2 so 12(-2) + 4y = -12 4y = -12 + 24 = 12 y = 12/4 = 3

OpenStudy (cwrw238):

now replace x by -2 and y by 3 in the equations in your question and see if they fit

OpenStudy (anonymous):

Wait wait.. How do you know x = -2?

OpenStudy (cwrw238):

look at my first post

OpenStudy (cwrw238):

15x = -30 so x = -2

OpenStudy (anonymous):

Oh okay i see..

OpenStudy (cwrw238):

so try the first equation in the question 6x + 2y = -6 x=-2, y=3 so we have 6(-2) + 2(3) = -12 + 6 = -6 so that one fits now try the same values in the second equation: 3x - 4y = -18

OpenStudy (cwrw238):

if they fit then you have your answer

OpenStudy (anonymous):

This is what you said: I'll find y for you in the first option: 12x + 4y = -12 now we found x = -2 so 12(-2) + 4y = -12 4y = -12 + 24 = 12 y = 12/4 = 3 So i understand that you plug in -2 into x so you would get 12(-2) + 4y = -12 But then in the next step i kinda got confused again.. 4y = -12 + 24 = 12 so you what did you exactly do? So would you do? Did you multiply 12 by -2?

OpenStudy (anonymous):

What did you do to get 4y all by itself on one side of the equation?

OpenStudy (cwrw238):

12(-2) + 4y = -12 -24 + 4y = -12 add 24 to both sides of the equation -24 + 24 + 4y = -12 + 24 4y = -12+24 = 12 now divide both sides by 4 y = 3

OpenStudy (solomonzelman):

\(\large\color{black}{ 6x + 2y = -6 \\ 3x - 4y = -18 }\) 2nd equation times -2. \(\large\color{black}{ 6x + 2y = -6 \\ -6x +8y = 36 }\) and the equations and tell me what you get for y.

OpenStudy (anonymous):

12(-2) + 4y = -12 -24 + 4y = -12 add 24 to both sides of the equation -24 + 24 + 4y = -12 + 24 4y = -12+24 = 12 now divide both sides by 4 y = 3 Okay i understand everything up until 4y = -12+24 = 12 I so understand how 4y = -12+24 got there but where did the = 12 come from?

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@cwrw238 will cross mutiplication method be helpful??

OpenStudy (anonymous):

Oh wait im so dumb... -12 + 24 = 12....

OpenStudy (anonymous):

So then you would be left with 4y = 12 then you said to divide both sides so x = 3

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

Okay thank you so much xD youre so helpful !!!!

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

Wait before you go.. I have one more question.... the equation 8x + 4y = -4 14x = -10 Would you divide 10 by 14 to find out what x equals?

OpenStudy (anonymous):

@SolomonZelman